OpenStudy (anonymous):

s=2πr^2+2πrh solve for h step by step please =]

OpenStudy (whynot):

s-2pi r^2=2pirh h=(s-2pi r^2)/2pi r

OpenStudy (anonymous):

I'll add some words. Given that s = 2*pi*r^2 + 2*pi*r*h subtract the term that does not contain h on the right hand side and add it to the left s - 2*pi*r^2 = 2*pi*r*h Divide by the 2*pi*r on both sides to get (s - 2*pi*r^2)/(2*pi*r) = h

OpenStudy (anonymous):

Or rather, "subtract the term that does not contain h on the right hand side and add it to the left" should be "subtract the term that does not contain h on the right hand side and subtract it from the left", sorry brain dead there